the market demand for a good is P = 90-Q. The good can be produced at a constant cost of $10. How much producer surplus is created if the market is served by a monopolist as opposed to a competitive market?
Answer #1

Answer

The monopolist produces at MR=MC

MR=90-2Q An MR curve is double sloped than an inverse linear demand curve.

MC=10

equating it

90-2Q=10

2Q=80

Q=40

P=90-40=50

PS =area below the demand curve and above price and left of the quantity so it is a rectangle area

=difference between MC and price * quantity

=(50-10)*40

=$1600

the producer surplus is $1600

The perfect competitive market has MC=P so there is no producer surplus so the producer surplus increases by $1600